# Holographic spectral function of fermion in instantonic plasma

### v1

@@ Abstract
Using the gauge-gravity duality, we investigate the fermionic correlation function in the D(-1)-D3 brane system which describes the instantonic plasma in holography. In this system, the charge of the D(-1) brane as the D-instanton gives the gluon condensate. To simplify the holographic setup, we first reduce briefly the ten-dimensional supergravity background produced by D(-1)-D3-branes to an equivalently five-dimensional background. Then using the standard method for computing the Green function in the AdS/CFT dictionary, we derive the equations for the fermionic correlation functions and solve them numerically with the infalling boundary condition. Our numerical results illustrate that the fermionic correlation function as spectral function includes two branches of the dispersion curves whose behavior is very close to the results obtained from the method of hard thermal loop. And the effective mass generated by the medium effect of fermion splits into two values due to the spin-dependent interactions induced by instantons. Therefore, this work in holography demonstrates the instantonic configuration is very influential to the fermionic plasmino in plasma.
@@ 1 Introduction
However, due to the property of asymptotic freedom, it is very challenging to investigate QCD in low-energy region by using the perturbative method in quantum field theory (QFT) since QCD is strongly coupled in low-energy region. Fortunately, the gauge-gravity duality and AdS/CFT correspondence provide an alternative way to study the strongly coupled QFT [14, 15, 16] in holography. In the top-down approach, instanton is constructed simply as a geometric background with instantonic D-branes (D-instanton) [17, 18, 19, 20, 21], for example, the black D(-1)-D3 brane system can describe holographically the plasma with instanton in which the D(-1)-brane plays the role of instanton, and its dual field theory is $\mathcal{N}=4$ super Yang-Mills theory with instanton or theta angle. So there have been many works based on the holographic instanton e.g. thermodynamics of Yang-Mills theory [22], hadronic spectrum [23, 24], real-time dynamics [25], Chern-Simons theory [26, 27], chiral symmetry breaking [28], quark potential [29, 30, 31], Schwinger effect [32, 33, 34], the hadronic interaction [35, 36]. Despite all this, the study of fermion with the instanton remains to be blank to the existing works in holography although it attracts great interests as it is reviewed in [1, 2, 3] in QFT. Therefore, the goal of this work is to explore the fermionic correlation function with instanton as a start point for the further continuous investigations, since it can be evaluated by using the most fundamental principle in holography.
@@
To our goal, the setup in this work starts with the D(-1)-D3 brane system since it describes simply the instantonic plasma in holography, then we reduce briefly the 10d IIB supergravity background produced by black D3-branes with D(-1)-branes [17, 25, 28] to an effectively five-dimensional (5d) geometric background. Afterwards, we probe a bulk fermion in the 5d geometric background and derive its covariant Dirac equation. In order to evaluate the fermionic correlation function in the dual theory, we employ the standard technique in the AdS/CFT dictionary so that obtain the differential equations for the fermionic correlators in the bulk. By imposing the infalling boundary condition at the horizon, we afterwards solve the fermionic correlation functions numerically. Our numerical calculation illustrates that, without the D-instanton, there is only one branch of the holographic dispersion curves for fermion lying on the light cone which is similar as the existing works about holographic fermions [35, 36, 37, 38, 39, 40]. In the presence of the D-instanton, the fermionic correlation function has a correction which in leading order splits the dispersion curves into two branches at small momentum, and this behavior is very close to the dispersion curves from the approach of hard thermal loop (HTL) in QFT. Furthermore, the two branches of the dispersion curves reveals that the effective bound mass of fermion in plasma has two values in the presence of the D-instanton. Remarkably it agrees qualitatively with effective potential for the spin-dependent interactions induced by instantons computed by the process of one-gluon exchange in QCD [1, 2]. Therefore, this work, as a start point, is a holographic reproduction of the properties of fermion and instanton and we believe the instantonic configuration in plasma is very influential to the fermionic plasmino.
@@
The outline of this manuscript is given as follows. In Section 2, we collect the essentials of the 10d supergravity background produced by D3-branes with D(-1)-branes as D-instantons, and reduce it briefly to an effectively 5d gravity background. In Section 3, we demonstrate the principle in AdS/CFT dictionary to compute the holographic correlation function for spinor. Then we derive the covariant Dirac equation and the holographic counter term for spinor in order to obtain the differential equations for the fermionic correlators. In Section 4, we solve numerically the holographic fermionic correlation functions order by order with physical analysis and interpretation. The final section is the summary.
@@
2 The holographic background for instantonic plasma
In this section, let us briefly review the holographic background for instantonic plasma which corresponds to the system consisted of $N_{c}$ D3-branes with $N_{\mathrm{D}}$ D-instantons i.e. the D(-1)-branes in the large-$N_{c}$ limit in the type IIB string theory [17, 18, 25]. Then, we integrate out the $S^{5}$ part in order to obtain a five-dimensional (5d) background geometry for simplification.
@@ 2.1 The D3-brane background with D-instanton
$\left\langle\mathrm{Tr}F_{\mu\nu}F^{\mu\nu}\right\rangle\simeq\frac{N_{\mathrm{D}}}{16\pi^{2}V_{4}}=\frac{Q}{\left(2\pi\alpha^{\prime}\right)^{2}R^{4}}\frac{N_{c}}{\left(2\pi\right)^{4}},\ \mu,\nu=0,1...3$ (2.4)
@@ 2.2 The effective 5d background
So the action (2.1) becomes,
@@
Accordingly, the 5d solution for the metric obtained from (2.2) satisfying the equations of motion by varying action (2.6) is
@@
$ds_{\left(5\mathrm{d}\right)}^{2}=$ $e^{\frac{\phi}{2}}\frac{R^{2}}{z^{2}}\left\{-f\left(z\right)dt^{2}+d\mathbf{x}\cdot d\mathbf{x}+\frac{dz^{2}}{f\left(z\right)}\right\},$
@@
in which the holographic boundary is located at $z\rightarrow 0$ to continue our discussion.
@@ 3 The fermionic correlation function
In this section, we consider a probe fermion in the bulk geometry and evaluate its two-point Green function in holography by employing the standard method in the AdS/CFT dictionary. In particular, we will focus on the influence of the instanton charge denoted by $q$ in the fermionic correlation function.
@@ 3.1 General setup for the fermionic correlation function
Let us start with the principle of the AdS/CFT which illustrates that the generating functional of the dual conformal field theory (CFT) $Z_{\mathrm{CFT}}$ is equal to its gravitational partition function $Z_{\mathrm{gravity}}$ in the bulk geometry. Namely, for a spinor field $\psi$ in the 5d bulk, we have
@@
$S_{\mathrm{gravity}}^{ren}$ $=\int_{\mathcal{M}}\mathcal{L}_{\mathrm{gravity}}^{ren}\left[\bar{\psi},\psi\right]\sqrt{-g}d^{5}x$ (3.2)

### v2

# Holographic spectroscopy of fermion with instantons

@@ Abstract
Using the gauge-gravity duality, we investigate the fermionic spectroscopy in the D(-1)-D3 brane system. The background geometry of this system described by IIB supergravity includes a black (deconfined) and bubble (confined) D3-brane which corresponds respectively to a deconfined and a confined gauge theory in holography. The charge of the D(-1) brane as the D-instanton gives the gluon condensate in this model. To simplify the holographic setup, we first reduce briefly the ten-dimensional supergravity background produced by D(-1)-D3-branes to an equivalently five-dimensional background. Then the fermionic spectrum in the confined case is obtained by decomposing the fermion with dimensional reduction. In addition, by using the standard method for computing the Green function in the AdS/CFT dictionary, we derive the equations for the fermionic correlation functions and solve them numerically with the infalling boundary condition. Our numerical results in the deconfined case illustrate that the fermionic correlation function as spectral function includes two branches of the dispersion curves whose behavior is very close to the results obtained from the method of hard thermal loop. And the effective mass generated by the medium effect of fermion splits into two values due to the spin-dependent interactions induced by instantons. In the confined case, the holographic correlation function indicates several separated dispersion curves which illustrates consistently the onset mass in the fermionic spectrum we obtained. Therefore, this work in holography demonstrates the instantonic configuration is very influential to the fermion in QCD.
@@ 1 Introduction
However, due to the property of asymptotic freedom, it is very challenging to investigate QCD in low-energy region by using the perturbative method in quantum field theory (QFT) since QCD is strongly coupled in low-energy region. Fortunately, the gauge-gravity duality and AdS/CFT correspondence provide an alternative way to study the strongly coupled QFT [14, 15, 16] in holography. In the top-down approach, instanton is constructed simply as a geometric background with instantonic D-branes (D-instanton) [17, 18, 19, 20, 21], for example, the black D(-1)-D3 brane system can describe holographically the plasma with instanton in which the D(-1)-brane plays the role of instanton, and its dual field theory is $\mathcal{N}=4$ super Yang-Mills theory with instanton or theta angle. So there have been many works based on the holographic instanton e.g. thermodynamics of Yang-Mills theory [22], hadronic spectrum [23, 24], real-time dynamics [25], Chern-Simons theory [26, 27], chiral symmetry breaking [28], quark potential [29, 30, 31], Schwinger effect [32, 33, 34], the hadronic interaction [35, 36]. Despite all this, the study of fermion with the instanton remains to be blank to the existing works in holography although it attracts great interests as it is reviewed in [1, 2, 3] in QFT. Therefore, the goal of this work is to explore the fermion with instanton as a start point for the further continuous investigations, since it can be evaluated by using the most fundamental principle in holography.
@@
To our goal, the setup in this work starts with the D(-1)-D3 brane system since it describes simply the instantons in gauge theory through gauge-gravity duality [17, 25, 28]. Then we reduce briefly the 10d IIB supergravity background produced by black D3-branes with D(-1)-branes to an effectively five-dimensional (5d) geometric background. We note that the supergravity solution for the D(-1)-D3 brane system includes a black (deconfined) and a bubble (confined) D3-brane solution which corresponds respectively to a deconfined and a confined gauge theories with instantons in holography [37, 38]. Afterwards, we probe a bulk fermion in the 5d geometric background and derive its covariant Dirac equation. In the confined geometry, we decompose the bulk fermion with dimensional reduction, evaluate numerically the holographic spectrum of fermion. Since the bulk fermion must be a gauge-invariant operator, we identify it to baryon in phenomenology. In addition, we investigate the fermionic two-point correlation function in the dual theory by employing the standard technique in the AdS/CFT dictionary so that obtain the differential equations for the fermionic correlators in the bulk. By imposing the infalling boundary condition at the horizon, we afterwards solve the fermionic correlation functions numerically both in deconfined and confined background. Our numerical calculation in the deconfined case illustrates that, without the D-instanton, there is only one branch of the holographic dispersion curves for fermion lying on the light cone which is similar as the existing works about holographic fermions [35, 36, 39, 40, 41, 42]. In the presence of the D-instanton, the fermionic correlation function has a correction which in leading order splits the dispersion curves into two branches at small momentum, and this behavior is very close to the dispersion curves from the approach of hard thermal loop (HTL) in QFT. Furthermore, the two branches of the dispersion curves reveals that the effective bound mass of fermion in plasma has two values in the presence of the D-instanton. Remarkably it agrees qualitatively with effective potential for the spin-dependent interactions induced by instantons computed by the process of one-gluon exchange in QCD [1, 2]. In the confined case, the holographic correlation function indicates several separated dispersion curves which agrees with the fermionic spectrum we obtained. Therefore, this work, as a start point, is a holographic reproduction of the properties of fermion and instanton, we believe the instantonic configuration in QCD is very influential to the fermion.
@@
The outline of this manuscript is given as follows. In Section 2, we collect the essentials of the 10d supergravity background produced by D3-branes with D(-1)-branes as D-instantons, and reduce it briefly to an effectively 5d gravity background. In Section 3, we decompose the bulk spinor with dimensional reduction, then evaluate the eigen mass of fermion numerically. In Section 4, we demonstrate the principle in AdS/CFT dictionary to compute the holographic correlation function for spinor. Then we derive the covariant Dirac equation and the holographic counter term for spinor in order to obtain the differential equations for the fermionic correlators. In Section 4, we solve numerically the holographic fermionic correlation functions order by order with physical analysis and interpretation, both in the deconfined and confined case. The final section is the summary.
@@
2 The instantonic background in holography
In this section, let us briefly review the instantonic background which corresponds to the system consisted of $N_{c}$ D3-branes with $N_{\mathrm{D}}$ D-instantons i.e. the D(-1)-branes in the large-$N_{c}$ limit in the type IIB string theory [17, 18, 25]. Then, we integrate out the $S^{5}$ part in order to obtain a five-dimensional (5d) background geometry for simplification.
@@ 2.1 The D3-brane background with D-instanton
$\left\langle\mathrm{Tr}F_{\mu\nu}F^{\mu\nu}\right\rangle\simeq\frac{N_{\mathrm{D}}}{16\pi^{2}V_{4}}=\frac{Q}{\left(2\pi\alpha^{\prime}\right)^{2}R^{4}}\frac{N_{c}}{\left(2\pi\right)^{4}},\ \mu,\nu=0,1...3.$ (2.4)
Besides, by following [43, 44], there is a confining background which can be obtained from (2.2). That is to compactify one of the spatial dimension $x^{3}=y$ of the D3-brane on a circle $S^{1}$, so the worldvolume theory on the D3-branes is effectively 3d below the energy scale $M_{KK}=\frac{2\pi}{\delta y}$, where $\delta y$ refers to the size of $S^{1}$. Then identify the bulk supergravity geometry that has this geometry for its boundary, which resultantly leads to a double Wick rotation ($t\rightarrow-iy,y\rightarrow-it$) on the background presented in (2.2) as,
$ds^{2}$ $=e^{\frac{\phi}{2}}\left\{\frac{r^{2}}{R^{2}}\left[\eta_{\mu\nu}dx^{\mu}dx^{\nu}+f\left(r\right)dy^{2}\right]+\frac{1}{f\left(r\right)}\frac{R^{2}}{r^{2}}dr^{2}+R^{2}d\Omega_{5}^{2}\right\},\mu,\nu=0,1,2.$
$f\left(r\right)$ $=1-\frac{r_{KK}^{4}}{r^{4}}.$ (2.5)
The background (2.5) is defined only for $r>r_{KK}$ and it does not have a horizon, so $r_{KK}$ is the end of the spacetime. Since the warp factor $e^{\frac{\phi}{2}}\frac{r^{2}}{R^{2}}$ never goes to zero, the Wilson loop asymptotics leads to an area law which exhibits as confinement in the dual theory.
@@ 2.2 The effective 5d background
So in Einstein frame, the action (2.1) becomes,
@@
Since the supergravity solution (2.2) implies the kinetic terms of $\Phi,C_{0}$ presented in (2.7) cancel each other, the 5d solution is nothing but $\mathrm{AdS_{5}}$. Thus, in string frame it is given as,
@@
$ds_{\left(5\mathrm{d}\right)}^{2}=$ $e^{\frac{\phi}{2}}\frac{R^{2}}{z^{2}}\left[-f\left(z\right)dt^{2}+d\mathbf{x}\cdot d\mathbf{x}+\frac{dz^{2}}{f\left(z\right)}\right],$
@@
in which the holographic boundary is located at $z\rightarrow 0$ to continue our discussion. Moreover, the confined geometry presented (2.5) also reduces to an effective 5d background as,
$ds_{\left(5\mathrm{d}\right)}^{2}=$ $e^{\frac{\phi}{2}}\frac{R^{2}}{z^{2}}\left[\eta_{\mu\nu}dx^{\mu}dx^{\nu}+f\left(r\right)dy^{2}+\frac{dz^{2}}{f\left(z\right)}\right],$
$f\left(z\right)=$ $1-\frac{z^{4}}{z_{KK}^{4}},e^{\phi}=1-q\ln f\left(z\right),$
$z=$ $\frac{R^{2}}{r},z_{H}=\frac{R^{2}}{r_{KK}},q=\frac{z_{KK}^{4}Q}{R^{8}}.$ (2.11)
3 The spectrum of the baryonic fermion
In this section, our concern is the fermionic spectrum in the confined geometry. As the Wilson loop asymptotics leads to an area law, we consider a fermionic operator in the boundary theory which describes a confinement state. In particular, the fermionic operator should be gauge-invariant according to gauge-gravity duality  [^4] As a bottom-up approach, we will not attempt to discuss the construction of a baryon in detail. How to construct a baryon state can be reviewed in [36] in which the top-down approaches are employed. [15, 16, 44]. As the confined metric given in (2.11) becomes 3d below the energy scale $M_{KK}$, so it reduces to the following metric,
$ds_{\left(4\mathrm{d}\right)}^{2}$ $=e^{\frac{\phi}{2}}\frac{R^{2}}{z^{2}}\left[\eta_{\mu\nu}dx^{\mu}dx^{\nu}+\frac{dz^{2}}{f\left(z\right)}\right],$
$\equiv g_{xx}\eta_{\mu\nu}dx^{\mu}dx^{\nu}+g_{zz}dz^{2}.$ (3.1)
We start with a bulk fermion $\psi$ as a dual operator to the baryonic fermion at the boundary propagating on the background (3.1). Its Dirac action is given as,
$S_{\mathrm{bulk}}=i\int d^{4}x\sqrt{-g}\bar{\psi}\left(\Gamma^{M}\nabla_{M}-m\right)\psi,$ (3.2)
where the Dirac operator is computed as,
$\Gamma^{M}\nabla_{M}=\frac{1}{\sqrt{g_{xx}}}\gamma^{\mu}\partial_{\mu}+\frac{1}{\sqrt{g_{zz}}}\gamma\partial_{z}+\frac{3}{4}\frac{\partial_{z}\ln g_{xx}}{\sqrt{g_{zz}}}\gamma.$ (3.3)
So the Dirac equation (3.2) can be rewritten as,
[…]
where the coefficients are
$A_{1}=\sqrt{-\frac{g}{g_{xx}}},A_{2}=\sqrt{-\frac{g}{g_{zz}}},A_{3}=\frac{3}{4}\partial_{z}\ln g_{xx}\sqrt{-\frac{g}{g_{zz}}},A_{4}=m\sqrt{-g}.$ (3.5)
Let us further decompose the spinor $\psi$ by a series of the basis function as,
$\psi=\sum_{n,k}\left(\begin{array}[]{c}\psi_{+}^{\left(n\right)}f_{+}^{\left(n\right)}\\ \psi_{-}^{\left(k\right)}f_{-}^{\left(k\right)}\end{array}\right),$ (3.6)
where $\psi_{\pm}^{\left(k\right)}$ depends on $x^{\mu}$ and $f_{\pm}^{\left(k\right)}$ depends on $z$. Then insert the decomposition to the action (3.4), it becomes
$S_{\mathrm{bulk}}=$ $-\sum_{k,n}\int d^{3}xdz\bigg{[}\psi_{-}^{\left(k\right)\dagger}f_{-}^{\left(k\right)}A_{2}\psi_{+}^{\left(n\right)}\partial_{z}f_{+}^{\left(n\right)}+\psi_{-}^{\left(k\right)\dagger}f_{-}^{\left(k\right)}A_{3}\psi_{+}^{\left(n\right)}f_{+}^{\left(n\right)}$
$+\psi_{-}^{\left(k\right)\dagger}f_{-}^{\left(k\right)}A_{4}\psi_{+}^{\left(n\right)}f_{+}^{\left(n\right)}+\psi_{-}^{\left(k\right)\dagger}f_{-}^{\left(k\right)}A_{1}i\sigma^{\mu}\partial_{\mu}\psi_{-}^{\left(n\right)}f_{-}^{\left(n\right)}$
$+\psi_{+}^{\left(n\right)\dagger}f_{+}^{\left(n\right)}A_{1}i\bar{\sigma}^{\mu}\partial_{\mu}\psi_{+}^{\left(k\right)}f_{+}^{\left(k\right)}-\psi_{+}^{\left(n\right)\dagger}f_{+}^{\left(n\right)}A_{2}\psi_{-}^{\left(k\right)}\partial_{z}f_{-}^{\left(k\right)}$
$-\psi_{+}^{\left(n\right)\dagger}f_{+}^{\left(n\right)}A_{3}\psi_{-}^{\left(k\right)}f_{-}^{\left(k\right)}+\psi_{+}^{\left(n\right)\dagger}f_{+}^{\left(n\right)}A_{4}\psi_{-}^{\left(k\right)}f_{-}^{\left(k\right)}\bigg{]}.$ (3.7)
By imposing the normalization condition for the basis function,
[…]
with eigen equations
$\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{F_{-}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}-A_{4}}{A_{1}}F_{-}^{\left(n\right)}$ $=-M_{n}F_{+}^{\left(n\right)},$
$\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{F_{+}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}+A_{4}}{A_{1}}F_{+}^{\left(n\right)}$ $=M_{n}F_{-}^{\left(n\right)},$ (3.9)
action (3.7) reduces to the standard form of kinetic term for fermion as,
$S=-\sum_{n}\int d^{3}x\bigg{\{}\psi_{-}^{\left(n\right)\dagger}i\sigma^{\mu}\partial_{\mu}\psi_{-}^{\left(n\right)}+\psi_{+}^{\left(n\right)\dagger}i\bar{\sigma}^{\mu}\partial_{\mu}\psi_{+}^{\left(n\right)}+M_{n}\left[\psi_{-}^{\left(n\right)\dagger}\psi_{+}^{\left(n\right)}+\psi_{+}^{\left(n\right)\dagger}\psi_{-}^{\left(n\right)}\right]\bigg{\}}.$ (3.10)
The eigenvalue $M_{n}$ can be evaluated by solving numerically the following eigen equations
$\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{A_{2}}{A_{1}}\partial_{z}\left(\frac{F_{+}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}+A_{4}}{A_{1}}F_{+}^{\left(n\right)}\right)+\frac{A_{3}-A_{4}}{A_{1}}\left(\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{F_{+}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}+A_{4}}{A_{1}}F_{+}^{\left(n\right)}\right)$
$=-M_{n}^{2}F_{+}^{\left(n\right)},$
$\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{A_{2}}{A_{1}}\partial_{z}\left(\frac{F_{-}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}-A_{4}}{A_{1}}F_{-}^{\left(n\right)}\right)+\frac{A_{3}+A_{4}}{A_{1}}\left(\frac{A_{2}}{\sqrt{A_{1}}}\partial_{z}\left(\frac{F_{-}^{\left(n\right)}}{\sqrt{A_{1}}}\right)+\frac{A_{3}-A_{4}}{A_{1}}F_{-}^{\left(n\right)}\right)$
$=-M_{n}^{2}F_{-}^{\left(n\right)}.$ (3.11)
We note that, the mass $m$ presented in (3.2) is the only parameter to fit the experimental data. For massless fermion $m=0$, the equations for $F_{\pm}^{\left(n\right)}$ are identical. For $m\neq 0$, the mass spectrum of $M_{n}^{\pm}$ is separated with the separation $\sim m/2$. According to (3.6), the “$\pm$” can be identified to the proper quantum numbers of a baryon e.g. the 3rd component of the isospin. In this sense, we solve (3.11) with sufficiently small mass parameter $m=-0.01$ and the associated mass spectrum is given in Table1
$q=0$ $n=1$ $n=2$ $n=3$ $n=4$
$M_{+}$ 1.7899 4.2721 6.7597 9.2568
$M_{-}$ 1.8131 4.2978 6.7868 9.2853
Table 1: The fermionic spectrum with $m=0$ in the unit of $M_{KK}=1$.
and its dependence on $q$ is illustrated in Figure. The numerical results regard to the fact that the mass of baryons as the up and down components of a isospin is roughly equal. The mass spectrum is less dependent on the instanton density. However the separation of the eigen mass does not match well to the experimental data for the lowest baryon mass. For example, if we identify the states of mass $M_{1}^{\pm},M_{2}^{\pm}$ as proton ($n=1,I=+1/2$), neutron ($n=1,I=-1/2$), $\Sigma^{+}$ ($n=1,I=+1/2$), $\Sigma^{-}$ ($n=1,I=-1/2$), the mass ratio of $\Sigma^{+}$ and proton given in this model is $M_{2}^{+}/M_{1}^{+}\simeq 2.39$ less close to the experimental data $M_{\mathrm{proton}}/M_{\Sigma^{+}}\simeq 1.27$. The possible reason may be that the current theory is 3d instead of a 4d theory.
Figure 1: The mass spectrum as a function of the instanton density $q$.
@@ 4 The fermionic correlation function
In this section, we focus on the fermionic correlation function as a parallel method to investigate the fermionic spectrum. By introducing a probe fermion in the bulk geometry, we will evaluate its two-point Green function in holography by employing the standard method in the AdS/CFT dictionary. Since the fermion must be gauge-invariant operator, it could be a baryonic fermion. In particular, the influence of the instanton charge denoted by $q$ in the fermionic correlation function is also concerned.
@@ 4.1 General setup for the fermionic correlation function
Let us start with the principle of the AdS/CFT which illustrates that the generating functional of the dual conformal field theory (CFT) $Z_{\mathrm{CFT}}$ is equal to its gravitational partition function $Z_{\mathrm{gravity}}$ in the bulk geometry. Namely, for a spinor field $\psi$ in the D+1 dimensional bulk, we have
@@
$S_{\mathrm{gravity}}^{ren}$ $=\int_{\mathcal{M}}\mathcal{L}_{\mathrm{gravity}}^{ren}\left[\bar{\psi},\psi\right]\sqrt{-g}d^{D+1}x$ (4.2)